f you want to make 5.62 L of 0.200 M potassium dichromate solution, what mass of potassium dichromate must you weigh out? Include your units.
Answer #1

volume , V = 5.62 L

use:

number of mol,

n = Molarity * Volume

= 0.2*5.62

= 1.124 mol

Molar mass of K2Cr2O7,

MM = 2*MM(K) + 2*MM(Cr) + 7*MM(O)

= 2*39.1 + 2*52.0 + 7*16.0

= 294.2 g/mol

use:

mass of K2Cr2O7,

m = number of mol * molar mass

= 1.124 mol * 2.942*10^2 g/mol

= 3.307*10^2 g

Answer: 331 g